OpenStudy (liv33):

HELP!! A bird drops a stick to the ground from a height of 80 ft. The function h=-16t^2+80 gives the stick's approximate height h above the ground, in feet, after t seconds. Graph the function. At about what time does the stick hit the ground?

Directrix (directrix):

h(x) = -16t^2 + 80 Stick touches the ground when h = 0.

Directrix (directrix):

Solve for t. After that, we'll draw the graph. -16t^2 + 80 = 0 @Liv33

OpenStudy (liv33):

ok 1 min

Directrix (directrix):

-16t^2 + 80 = 0 -16t^2 = - 80 Then, what comes next?